OpenStudy (anonymous):

if f(x)=-6x+6, then f^-1(x)

OpenStudy (usukidoll):

any attempts?

OpenStudy (anonymous):

solve \[x=-6y+6\] for \(y\) in two steps

OpenStudy (usukidoll):

that's what I was thinking.. swap the variables.. solve for y and then replace y with the f inverse notation .

OpenStudy (anonymous):

at least that is the math teacher way the thinking way is to recognize that\[f(x)=-6x+6\] means a) multiply by \(-6\) then b) add \(6\) inverse will do the opposite operations in reverse order

OpenStudy (anonymous):

but solving \[x=-6y+6\] for \(y\) takes two steps, so it is just as quick as thinking

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

can you work through it with me?

OpenStudy (anonymous):

please? all my attempts arent working out

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (anonymous):

Just follow these easy steps: To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

I know the steps but I usually get stuck when solving for y

OpenStudy (anonymous):

\[y=-6x+6 \] \[x=-6y+6\] what would your first step be to solving for y?

OpenStudy (anonymous):

Try your best, don't be afraid to make mistakes

OpenStudy (anonymous):

get rid of + 6 by subtracting six from both sides

OpenStudy (anonymous):

There you go, good job, so we have \[x-6 = -6y\] now what?

OpenStudy (anonymous):

divide by negative 6

OpenStudy (anonymous):

Good! \[y = \frac{ -(x-6) }{ 6 }\] you may simplify that now

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

As in you can distribute the negative sign, and get \[y = \frac{ -x+6 }{ 6 } \implies \frac{ 6-x }{ 6 }\] so what's our final step?

OpenStudy (anonymous):

What should we replace y with :)

OpenStudy (anonymous):

the rule for negative functions

OpenStudy (anonymous):



OpenStudy (anonymous):

\[f^{-1}(x) = \frac{ 6-x }{ 6 }\] yup :)

OpenStudy (anonymous):

Not so bad right? Just have to try it and you'll start getting it!

OpenStudy (anonymous):

it helps to talk it out, I probably would've taken a lot longer to get that on my own

OpenStudy (anonymous):

I think that is the best way of learning to :)

OpenStudy (anonymous):

thanks for your help, i understand it better, not completely yet but making progress

OpenStudy (anonymous):

No problem, take care!